Proposal:
K Kids
Target of Oportunity:
No
Time Observed:
9.0
Time Engineering:
0.0
Time lost technical:
0.0
Time lost weather:
0.0
Time lost other:
0.0
Primary Investigator:
Todd Henry
Email:
Institution:
SMARTS
Instrument:
CTIO 1.5m Echelle
Organization:
SMARTS
Total Program Time:
9

- NOS line 35 was SENT however, status did not change.
- Air conditioning unit working at 10C (should be 16C), "ramps" to 35C, manually set back to 16C, goes down to 10C. 4UT ramps again to 35C, manually set back to 16C. I heard a noise, might be due to the cycle change (cooling to heating and back to cooling.) It is like a 1 hour 40 min cycle. End of night: turn it off and installed a fan to ventilate the room. The other noise I reported last night (email) was just the drivers relays.
- No time for VXR70 and PMM5884.
- Last object, HIP042507, ended with the Sun at -8.0 degrees.